OpenStudy (anonymous):

Find the value of k so that the line passing through the given points has slope m. Write an equation of the line in point-slope form. (k, 4k), (k+2,3k), m = -1

OpenStudy (anonymous):

given (k, 4k), (k+2,3k), m = -1 m=(4k-3k)/(k-(k+2)) m=1k/(-2)=-1 continue here and find k=? now using the point (k, 4k)=(?,?) using the value of k, and using m=-1 set up an Equation using the formula y - y1= m (x - x1) y - ? = -1 (x - ?) then simplify this to make point slope form

OpenStudy (anonymous):

Would it be y+2 = 1/-2 (x - .5) ?

OpenStudy (anonymous):

using the points and slope given (k, 4k), (k+2,3k), m = -1 (4k-3k) m= ---------- (k-(k+2)) 1k m = ----- = -1 -2 therefore k=2 now the points are (k, 4k)=(2,8) why? (k+2,3k)=(4,6) why? now using either of these points (2,8) or (4,6) and using the formula of point-slope form. y - y1= m( x - x1) y - 8 = -1(x - 2 ) yes or no?

OpenStudy (anonymous):

Because k=2 and it's just k in the first x and 4 times 2 is 8 so it's (2,8) and 2+2 = 4 and 3 times 2 = 6 so it's (4,6). Yes that's correct! Thank you! I just wasn't sure how you would find it. I kept getting 1 over -2. I just couldn't figure out what to do with it.